OpenStudy (anonymous):

5x - 2y for x = -7 and y = - 3

OpenStudy (anonymous):

5(-7)-2(-3) -35+6 -29

OpenStudy (anonymous):

-29

OpenStudy (anonymous):

-29

OpenStudy (anonymous):

@jdillon do u have a calculator??

OpenStudy (anonymous):

5(-7) - 2(-3) 5*-7 = -35 2*-3 = -6 -35 - -6 = -35 + 6 = -29